`I have a problem to solve it is
3x+2y=4
5x-4y=3
A. (1,0.5)
B. (0.5,1)
C. (5,5.5)
D. (5/11, -2/11)
I have no idea how to solve this and I tried the subratction method that wasnt
right so I know that I am doing something wrong can you help me
Thanks''

Answer by IBRAHIM92(5)

3x+2y=4
5x-4x=3
multiply the first equation x 2
6x+4y=8
Add the two equations
11x =11
x=1
hence y =1/2 or 0.5
answer = 1 ,0.5